OpenStudy (anonymous):

How do i prove: sin(theta)+sin(2theta)/cos(2theta)+cos(theta)+1=tan(theta)

OpenStudy (raden):

use these identities : sin(2θ) = 2sin(θ)cos(θ) and cos(2θ) = 2cos^2 (θ) - 1

OpenStudy (anonymous):

okayy. i believe i got it!

OpenStudy (raden):

so, (sin(θ)+sin(2θ))/(cos(2θ)+cos(θ)+1) = (sin(θ) + 2sinθcosθ)/(2cos^2 θ - 1 + cosθ+1) = (sin(θ) + 2sinθcosθ)/(2cos^2 θ + cosθ) factor out the numerator and the denominator by take the commons, get (sin(θ) + 2sinθcosθ)/(2cos^2 θ + cosθ) = sin(θ) (1+ cosθ)/(cosθ(2cosθ+1)) cancel the (1+cosθ)'s, get = sin(θ)/cos(θ) = tan(θ)

OpenStudy (raden):

i meant the last 3 is sin(θ) (1+ 2cosθ)/(cosθ(2cosθ+1)) so, cancel the (1+ 2cosθ)

OpenStudy (anonymous):

okay. great! thats what i got! thank you. i was wondering if you could also help me with.... prove: sec x - sin x cot x=sin^2x/cos x

OpenStudy (anonymous):

i know we need to change the sec x into (1/cos x) and cot x into (cos x/sin x) but after that, im not sure what to do...

OpenStudy (raden):

yeah, that's good starting to prove it.

OpenStudy (raden):

sec x - sin x cot x = 1/cosx - sin * cosx/sinx (cancel the sin's) = 1/cosx - cosx ?

OpenStudy (anonymous):

that is what i get but how do we get that to sin^2 x?

OpenStudy (raden):

1/cosx - cosx = 1/cosx - (cosx * cosx/cosx) = 1/cosx - cos^2 (x)/cosx see the denominator's already be same, so just adding the numerator. get = 1-cos^2 (x) / cosx

OpenStudy (raden):

finally, are you familiar with this identity : sin^2 (x) + cos^2 (x) = 1 ?

OpenStudy (anonymous):

= 1/cosx - (cosx * cosx/cosx) can you explain this?

OpenStudy (anonymous):

and yes i am.

OpenStudy (raden):

i just want the cos(x) be a fraction, it just a trick, look that cos(x) = cos(x)/1 if you times the numerator and the denominator by cos(x), get cos(x) = cos(x)/1 * cosx/cosx = cos(x)/cos^2 x, right ?

OpenStudy (raden):

damn,.. that's backward, i meant cos^2 x/cosx

OpenStudy (anonymous):

okay. i understand.

OpenStudy (anonymous):

thank you so much! i really appreciate it!

OpenStudy (raden):

you're welcome :)